mathslover (mathslover):

when two bulbs are connected individually of 110 volt then there are powers are 25 watt and 100 watt . If they are connected in series which will be more brighter ?

OpenStudy (anonymous):

Alright so you have 2 bulbs of 110 volts and one with 25 watt and another 100watt

OpenStudy (anonymous):

?

mathslover (mathslover):

right

OpenStudy (anonymous):

do you have a formula i can see? i've never seen a question like this

mathslover (mathslover):

no thts why m confused

OpenStudy (anonymous):

when they are in series the one wich has more resistence will bright more,since they have same current.\[P=r.i^{2}\]

OpenStudy (anonymous):

so it's R1

OpenStudy (anonymous):

^ You have to use Ohm's law and Watt's Law

OpenStudy (anonymous):

110²/R1=25 110²/R2=100 (R2/R1)=1/4 →R1=4.R2

mathslover (mathslover):

thanks very much Raphael Filgueiras

OpenStudy (anonymous):

@mathslover just replace is110^2/R1

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

P=V*I and V = IR So the current through the 100 wat is 100/110= 0.909 A, for the 25 watt; 25/110= 0.227 A. The resistance through the 100W = 110/.909=121 Ohms, for the 25W =110/.227=484 Ohms. So, in series, the total resistance of the circuit is 121+484=605 Ohms. The current in the circuit is 110/605= 0.181A. Multiply this by the voltage and you get; 110 * 0.181 = 20W. So they both are at the same brightness as that is the total current of the circuit. You can't get any more power ( "its the dylithium crystals, Captain"). This means the 25W will be dimmer than earlier. I think I'm correct here, unless I've ballsed up.

OpenStudy (anonymous):

PS That is why fairly lights are connected in parallel on your holiday tree.